2. Kapoor Denims Inc. manufactures denim fabric. 20 metres of the fabric is sold at Rs 8000. Fixed costs

is Rs 20 lacs per production period and the profit contribution is 40 percent of price. (20 marks)

a. Determine the BEP.

b. Determine the Profit/loss at output of 8,000, 10,000 and 15,000 units.

c. For the next production period, fixed costs will increase to 30 lacs due to a major capital

investment programme, but the new and more efficient machinery will result in a lower variable

production cost so that variable cost per metre will be reduced by 40 percent. If price is unchanged, re-

compute the profit /loss at output rate of 8,000, 10,000 and 15,000 units


1
Expert's answer
2021-12-03T11:50:26-0500

(a)

"BEP=\frac{Fixed Cost}{Contribution Margin}"

"=\frac{20}{0.40\times 8000}=0.00625"


(b)

Profit = Total Revenue- Total Cost

8,000:

"\frac{8000}{20}=400"

"=(400\times8000)-2000000"

"=1200000"


10,000:

"\frac{10000}{20}=500"

"=(500\times8000)-2000000"

"=2000000"


15,000:

"\frac{15000}{20}=750"

"=(750\times8000)-2000000"

"=4000000"


(c)

8,000:

"=(400\times8000)-3000000"

"=200000"


10,000:

"=(500\times8000)-3000000"

"=1000000"


15,000:

"=(750\times8000)-3000000"

"=3000000"